# Failure of famous functional inequalities on Finsler manifolds: the influence of $S$-curvature

### v1

@@ 4.1. Vanishing first eigenvalue of $p$-Laplacian
In this subsection we will study the first eigenvalue of the $p$-Laplacian in the Finsler setting. According to Shen [27], the divergence of a smooth vector field $X$ with respect to $\operatorname{\mathfrak{m}}$ is defined as
@@
where $r(x):=d_{F}(o,x)$. Then for any $p\in[1,\infty)$, one has
@@
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$
@@
$\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)\leq\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$
For the reverse inequality, it suffices to show that for any $u\in\mathsf{Lip}_{0}(M)\backslash\{0\}$, there exists a sequence $u_{l}\in C^{\infty}_{0}(M)\backslash\{0\}$ such that
@@
$\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}=\|{\operatorname{d}}u\|^{p}_{L^{p}_{\operatorname{\mathfrak{m}}}}=\lim_{n\rightarrow\infty}\|{\operatorname{d}}u_{l}\|^{p}_{L^{p}_{\operatorname{\mathfrak{m}}}}=\lim_{l\rightarrow\infty}\int_{M}F^{*p}({\operatorname{d}}u_{l})\operatorname{d\mathfrak{m}}.$ (4.14)
Thus, (4.9) follows directly by (4.11) and (4.14).
@@
If $\sup_{x\in M}r(x)<+\infty$, then $M$ is compact and $u_{\alpha}\in C^{\infty}(M)=C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$. Hence, (4.15) is trivial by choosing $v_{\alpha,l}=u_{\alpha}$. In the sequel, we consider the case when $\sup_{x\in M}r(x)=+\infty$, i.e., $M$ is noncompact. Thus, for every small $\varepsilon\in(0,\min\{1,\mathfrak{i}_{o}\})$ and $\delta\in(0,e^{-\alpha})$, set
@@
[…]
@@
thus, the limit in (4.16) follows.
@@
In view of (4.6) and (4.7), the first eigenvalue of the $p$($>1$)-Laplace operator $\Delta_{p,\operatorname{\mathfrak{m}}}$ is defined as
@@
for some constants $h\geq k\geq 0$. Thus $\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)=0$ for all $p\in(1,\infty)$.
@@
(i) Consider a noncompact Riemannian manifold equipped with the canonical Riemannian measure satisfying $\mathbf{Ric}\geq 0$. Corollary 4.1 implies the first eigenvalue of the Laplacian vanishes, which has been observed first by Cheng and Yau [6, Proposition 9].
@@
Let $(M,F)$ be an $n$-dimensional Cartan–Hadamard manifold with constant flag curvature $-k^{2}<0$. For each $p\in(1,\infty)$, the following statements hold::::
@@
$\lambda_{1,p,\operatorname{\mathfrak{m}}_{\rm can}}(\mathbb{H}^{n})=\frac{(n-1)^{2p}k^{2p}(p-1)^{p}}{p^{2p}}>0;$
see McKean [20] for $p=2$, and Farkas, Kajántó and Kristály [8] for generic $p>1$. Clearly, $\mathbf{S}=0$, which is consistent with Corollary 4.2/(b).
@@ 4.2. Failure of Hardy inequality
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u).$
@@
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)=0,$
@@
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u).$
@@
where ${\operatorname{d}}\mu=r^{-p}\operatorname{d\mathfrak{m}}$. Thus (4.24) follows because (4.14) remains valid.
@@
It is enough to discuss the case $\sup_{x\in M}r(x)=+\infty$; otherwise, $M$ is compact and (4.26) is obvious by choosing $v_{\alpha,l}=u_{\alpha}\in C^{\infty}_{0}(M)$. Thus, for every small $\varepsilon>0$, set
@@
Therefore, $\tilde{u}_{\alpha,\varepsilon}\in\mathsf{Lip}_{0}(M)$ follows by the same argument as in Step 2 of the proof of Lemma 4.2. Moreover, the dominated convergence theorem combined with (4.17) and Lemma 4.1/(i)&(ii) yields
@@
where $r(x):=d_{F}(o,x)$. Thus, the $L^{p}$-Hardy inequality fails for all $p\in(1,n)$, i.e.,
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}=0.$ (4.27)
@@
Case 1: $k>0$. Thus, (4.21) together with (4.23), (4.29) and (4.1) yields
@@
Thus, (4.23) together with (4.29) and (4.30) implies
@@
(i) A closer inspection of the proof of Theorem 4.2 clearly shows that if $h=0$, the last limit would be +$\infty$+ ∞ since $p<n$. This is not surprising since in the Riemannian setting, the $L^{p}$-Hardy inequality is valid on every $n$-dimensional Cartan–Hadamard manifold equipped with the canonical Riemannian measure, due to the vanishing $S$-curvature (2.12), see Yang, Su and Kong [34] and Kristály [16]. On the other hand, there are several Finsler manifolds satisfying the assumptions of Corollary 4.3; see Section 6 for details.
@@ 4.3. Failure of Heisenberg–Pauli–Weyl uncertainty principle
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\left(\displaystyle\int_{M}F^{*}({\operatorname{d}}u)^{2}\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}u^{2}r^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}=0.$ (4.31)
@@
where $u_{\alpha}:=-e^{-\alpha r}$ for $\alpha>0$. Thus it suffices to show $\lim_{\alpha\rightarrow 0^{+}}\mathcal{H}_{o}(u_{\alpha})=0$.
@@
$e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r)r^{2}<1\quad\text{for }r\in(\delta,\infty).$ (4.34)
@@
On the other hand, since (4.28) is always valid, we obtain
@@
(i) One cannot relax the assumption $h>k\geq 0$ to $h=0$ (and $k=0$). Indeed, according to Huang, Kristály and Zhao [10, Theorem 1.2], if a Cartan–Hadamard manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$ satisfies
@@
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{H}_{o}(u)\geq\frac{1}{\lambda_{F}(M)^{2}}\frac{n^{2}}{4}.$
@@
Indeed, due to Wu and Xin [33, Lemma 3.3 & Theorem 4.1], we have $\operatorname{\mathfrak{m}}$-a.e. in $M$ that
@@
The interpretation of the classical Heisenberg–Pauli–Weyl uncertainty principle in quantum mechanics is that the position and momentum of a given particle cannot be accurately determined simultaneously; this phenomenon can be captured in the presence of the ’gap’ $n^{2}/4$ in relation (1.3) (or the value $1/4$ in the non-sharp inequality (4.37)). However, in our geometric context provided by Theorem 4.3, the theoretical possibility of simultaneous detection of position and momentum of a particle still persist due to the gap-vanishing fact.
@@ 4.4. Failure of Caffarelli–Kohn–Nirenberg interpolation inequality
$\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\left(\displaystyle\int_{M}F^{*2}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}=0,$ (4.38)
@@
On the other hand, since $\eta:=h-k>0$, a similar argument yields $\delta=\delta(n,q,k,h)\in(\varepsilon,\infty)$ with the property that
$e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{(n-1)}(r)r^{2-2q}\leq e^{-\frac{(n-1)\eta}{2}r},\quad\forall\,r\in(\delta,\infty).$ (4.41)
@@
$\leq$ $\left(\int_{S_{o}M}e^{-\tau(y)}{\operatorname{d}}\nu_{o}(y)\right)\left(\int_{0}^{i_{y}}\left[1+(\alpha\,r)^{2-q}\right]^{\frac{2p-2}{2-p}}r^{2-2q}e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
@@
$=$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\int^{\varepsilon}_{0}+\int_{\varepsilon}^{\delta}+\int_{\delta}^{\infty}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(2\int^{\varepsilon}_{0}r^{n-2q+1}{\operatorname{d}}r+\int_{\varepsilon}^{\delta}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r+\int_{\delta}^{\infty}e^{-\frac{(n-1)\eta}{2}r}{\operatorname{d}}r\right)$
@@
=:absent:=:= : $C_{1}(n,q,k,h,\delta,\varepsilon)<+\infty.$ (4.42)
@@
As before, it is enough to consider the case when $\sup_{x\in M}r(x)=+\infty$; otherwise, $M$ is compact and the sequence would be $u_{\alpha,n}:=u_{\alpha}\in C^{\infty}(M)=C^{\infty}_{0}(M)\subset\mathsf{Lip}_{0}(M)$. Thus, for small $\varepsilon\in(0,1)$ set
@@
On can easily show that $v_{\alpha,\varepsilon},u_{\alpha,l}\in\mathsf{Lip}_{0}(M)$ since $u_{\alpha}\in C^{\infty}(M)$ with $\lim_{r(x)\rightarrow\infty}u_{\alpha}(x)=0$. In addition, the construction shows the pointwise convergences
@@
Thus (4.44) can be derived by these limits, (4.39) and the dominated convergence theorem.
@@
$=$ $\liminf_{\alpha\rightarrow 0^{+}}\alpha^{2(2-q)}\left(\frac{2-q}{p-2}\right)^{2}\left(\frac{C_{1}}{C_{2}}\right)^{2}=0,$
@@
5. Influence of $S$-curvature on geometric aspects: proof of Theorem 1.2

### v2

@@ 4.1. Vanishing first eigenvalue of $p$-Laplacian
In this subsection we will study the first eigenvalue of the $p$-Laplacian in the Finsler setting. According to [28], the divergence of a smooth vector field $X$ with respect to $\operatorname{\mathfrak{m}}$ is defined as
@@
with $r(x):=d_{F}(o,x)$. Then for any $p\in[1,\infty)$,
@@
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$
@@
$inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\Lambda_{p}(u)\leq\inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\Lambda_{p}(u).$
[…]
@@
$\int_{M}F^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}=\|{\operatorname{d}}u\|^{p}_{L^{p}_{\operatorname{\mathfrak{m}}}}=\lim_{l\rightarrow\infty}\|{\operatorname{d}}u_{l}\|^{p}_{L^{p}_{\operatorname{\mathfrak{m}}}}=\lim_{l\rightarrow\infty}\int_{M}F^{*p}({\operatorname{d}}u_{l})\operatorname{d\mathfrak{m}}.$ (4.14)
Then (4.9) follows directly by (4.11) and (4.14).
@@
If $\sup_{x\in M}r(x)<+\infty$, then $M$ is compact and $u_{\alpha}\in\mathsf{Lip}_{0}(M)$. Hence, (4.15) is trivial by choosing $v_{\alpha,l}=u_{\alpha}$. In the sequel, we consider the case when $\sup_{x\in M}r(x)=+\infty$, i.e., $M$ is noncompact. For every small $\varepsilon\in(0,\min\{1,\mathfrak{i}_{o}\})$ and $\delta\in(0,e^{-\alpha})$, set
@@
$|\tilde{u}_{\alpha,\varepsilon}|\nearrow|u_{\alpha}|,\quad F^{*}({\operatorname{d}}\tilde{u}_{\alpha,\varepsilon})\nearrow F^{*}({\operatorname{d}}u_{\alpha}),\quad\text{ for $\operatorname{\mathfrak{m}}$-a.e.,}$ (4.17)
@@
then the limit in (4.16) follows.
@@
In view of (4.6) and (4.7), the first eigenvalue of the pp(>1>1)-Laplace operator $\Delta_{p,\operatorname{\mathfrak{m}}}$ is defined as
@@
for some constants $h\geq k\geq 0$. Then $\lambda_{1,p,\operatorname{\mathfrak{m}}}(M)=0$ for all $p\in(1,\infty)$.
@@
(i) Consider a noncompact Riemannian manifold equipped with the canonical Riemannian measure satisfying $\mathbf{Ric}\geq 0$. Corollary 4.1 implies the first eigenvalue of the Laplacian vanishes, which has been observed first by Cheng–Yau [6, Proposition 9].
@@
Let $(M,F)$ be an $n$-dimensional Cartan–Hadamard manifold with constant flag curvature $k^{2}<0$. For each $p\in(1,\infty)$, the following statements hold::
@@
$\lambda_{1,p,\operatorname{\mathfrak{m}}_{\rm can}}(\mathbb{H}^{n})=\frac{(n-1)^{2p}k^{2p}(p-1)^{p}}{p^{2p}}>0,$
see McKean [20] for $p=2$, and Farkas–Kajántó–Kristály [8] for generic $p>1$. The fact that $\mathbf{S}=0$ in this case aligns perfectly with Corollary 4.2/(b).
@@ 4.2. Failure of Hardy inequality
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u).$
@@
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)=0,$
@@
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u)=\inf_{u\in\mathsf{Lip}_{0}(M)\backslash\{0\}}\mathcal{J}_{o,p}(u).$
@@
where ${\operatorname{d}}\mu=r^{-p}\operatorname{d\mathfrak{m}}$. Then (4.24) follows because (4.14) remains valid.
@@
It is enough to discuss the case $\sup_{x\in M}r(x)=+\infty$; otherwise, $M$ is compact and (4.26) is obvious by choosing $v_{\alpha,l}=u_{\alpha}\in\mathsf{Lip}_{0}(M)$. For every small $\varepsilon>0$, set
@@
Then $\tilde{u}_{\alpha,\varepsilon}\in\mathsf{Lip}_{0}(M)$ follows by the same argument as in Step 2 of the proof of Lemma 4.2. Moreover, the dominated convergence theorem combined with (4.17) and Lemma 4.1/(i)(ii) yields
@@
where $r(x):=d_{F}(o,x)$. Then the $L^{p}$-Hardy inequality fails for all $p\in(1,n)$, i.e.,
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\displaystyle\int_{M}{F}^{*p}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}}{\displaystyle\int_{M}\frac{|u|^{p}}{r^{p}}\operatorname{d\mathfrak{m}}}=0.$ (4.27)
@@
Case 1: $k>0$. For this case, combining (4.21), (4.23), (4.29), and (4.1) yields
@@
Then (4.23) together with (4.29) and (4.30) implies
@@
(i) A closer inspection of the proof of Theorem 4.2 clearly shows that if $h=0$, the last limit would be $\infty$ since $p<n$. This is not surprising since in the Riemannian setting, the $L^{p}$-Hardy inequality is valid on every $n$-dimensional Cartan–Hadamard manifold equipped with the canonical Riemannian measure, due to the vanishing $S$-curvature (1.1), see Yang–Su–Kong [35] and Kristály [16]. On the other hand, there are several Finsler manifolds satisfying the assumptions of Corollary 4.3; see Section 6 for details.
@@ 4.3. Failure of Heisenberg–Pauli–Weyl uncertainty principle
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\left(\displaystyle\int_{M}F^{*}({\operatorname{d}}u)^{2}\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}u^{2}r^{2}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}u^{2}\operatorname{d\mathfrak{m}}\right)^{2}}=0.$ (4.31)
@@
where $u_{\alpha}:=-e^{-\alpha r}$ for $\alpha>0$. Then it suffices to show $\lim_{\alpha\rightarrow 0^{+}}\mathcal{H}_{o}(u_{\alpha})=0$.
@@
$e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r)r^{2}<1\quad\text{for }r\in(\delta,+\infty).$ (4.34)
@@
On the other hand, since (4.28) is valid, we obtain
@@
(i) One cannot relax the assumption $h>k\geq 0$ to $h=0$ (and $k=0$). Indeed, according to [10, Theorem 1.2], if a Cartan–Hadamard manifold $(M,F)$ endowed with a smooth positive measure $\operatorname{\mathfrak{m}}$ satisfies
@@
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\mathcal{H}_{o}(u)\geq\frac{1}{\lambda_{F}(M)^{2}}\frac{n^{2}}{4}.$
@@
Indeed, due to Wu–Xin [34, Lemma 3.3, Theorem 4.1], we have $\operatorname{\mathfrak{m}}$-a.e. in $M$ that
@@
The interpretation of the classical Heisenberg–Pauli–Weyl uncertainty principle in quantum mechanics is that the position and momentum of a given particle cannot be accurately determined simultaneously; this phenomenon can be captured in the presence of the ‘gap’ $n^{2}/4$ in relation (1.4) (or the value $1/4$ in the non-sharp inequality (4.37)). However, in our geometric context provided by Theorem 4.3, the theoretical possibility of simultaneous detection of position and momentum of a particle still persist due to the gap-vanishing fact.
@@ 4.4. Failure of Caffarelli–Kohn–Nirenberg interpolation inequality
$inf_{u\in C^{\infty}_{0}(M)\backslash\{0\}}\frac{\left(\displaystyle\int_{M}F^{*2}({\operatorname{d}}u)\operatorname{d\mathfrak{m}}\right)\left(\displaystyle\int_{M}\frac{|u|^{2p-2}}{r^{2q-2}}\operatorname{d\mathfrak{m}}\right)}{\left(\displaystyle\int_{M}\frac{|u|^{p}}{r^{q}}\operatorname{d\mathfrak{m}}\right)^{2}}=0,$ (4.38)
@@
On the other hand, since $\eta:=h-k>0$, a similar argument yields $\delta=\delta(n,q,k,h)\in(\varepsilon,+\infty)$ with the property that
$e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{(n-1)}(r)r^{2-2q}\leq e^{-\frac{(n-1)\eta}{2}r},\quad\forall\,r\in(\delta,+\infty).$ (4.41)
@@
$\leq$ $\left(\int_{S_{o}M}e^{-\tau(y)}{\operatorname{d}}\nu_{o}(y)\right)\left(\int_{0}^{+\infty}\left[1+(\alpha\,r)^{2-q}\right]^{\frac{2p-2}{2-p}}r^{2-2q}e^{-(n-1)hr}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
@@
$=$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(\int^{\varepsilon}_{0}+\int_{\varepsilon}^{\delta}+\int_{\delta}^{+\infty}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r\right)$
$\leq$ $\mathscr{I}_{\operatorname{\mathfrak{m}}}(o)\left(2\int^{\varepsilon}_{0}r^{n-2q+1}{\operatorname{d}}r+\int_{\varepsilon}^{\delta}e^{-(n-1)hr}r^{2-2q}\mathfrak{s}_{-k^{2}}^{n-1}(r){\operatorname{d}}r+\int_{\delta}^{+\infty}e^{-\frac{(n-1)\eta}{2}r}{\operatorname{d}}r\right)$
@@
$=:$ $C_{1}(n,q,k,h,\delta,\varepsilon)<+\infty.$ (4.42)
@@
As before, it is enough to consider the case when $\sup_{x\in M}r(x)=+\infty$; otherwise, $M$ is compact and the sequence would be $u_{\alpha,n}:=u_{\alpha}\in\mathsf{Lip}_{0}(M)$. Then for small $\varepsilon\in(0,1)$, set
@@
On can easily show that $v_{\alpha,\varepsilon},u_{\alpha,l}\in\mathsf{Lip}_{0}(M)$ since $u_{\alpha}\in C^{\infty}(M)$ with $\lim_{r(x)\rightarrow+\infty}u_{\alpha}(x)=0$. In addition, the construction shows the pointwise convergences
@@
Then (4.44) can be derived by these limits, (4.39) and the dominated convergence theorem.
@@
$\leq$ $\liminf_{\alpha\rightarrow 0^{+}}\alpha^{2(2-q)}\left(\frac{2-q}{p-2}\right)^{2}\left(\frac{C_{1}}{C_{2}}\right)^{2}=0,$
@@
5. Influence of $S$-curvature on geometric aspects: proof of Theorems 1.2 &1.3